# Neuronal Attention Circuit (NAC) for Representation Learning

### v2

@@ A.1 Attention Mechanism
    Normalize the logits to get attention weights and compute the output:
@@ Appendix B Proofs
B.1 Deriving Closed-form (Exact) Solution
Although $\phi$ and $\omega_{\tau}$ are nonlinear functions of the input $\mathbf{u}=[\mathbf{q};\mathbf{k}]$, we derive closed-form solution by treating them as locally constant over the pseudo-time integration interval for each query–key pair based on frozen-coefficient approximation (john1952integration). This is accurate whenever the interval is short or when input variations are slow compared with the relaxation rate $\omega_{\tau}$. Under approximation assumption, rewrite Eqn. 1 as
@@ B.1 Deriving Closed-form (Exact) Solution
This is now a linear first-order ODE. The integrating factor is
@@
Recognize the left-hand side as the derivative of $e^{\omega_{\tau}t}a_{t}$:
@@
Compute the integral (since $\omega_{\tau}\neq 0$):
@@
Set $a^{*}:=\dfrac{\phi}{\omega_{\tau}}$. Then $a_{t}=a^{*}+(a_{0}-a^{*})e^{-\omega_{\tau}t}$, proved.
@@ C.1.1 Trajectory sensitivities for Closed-form formulation
C.2 Gradient-Based Training
@@
C.3 Efficiency and Complexity
Table 3 summarizes the computational complexity of different sequence models. For sequence prediction over length $n$ with hidden dimension $k$, RNNs scale linearly, $\mathcal{O}(nk)$, while Attention and NAC scale quadratically, $\mathcal{O}(n^{2}k)$. ODE-based models, such as LNNs, incur an additional multiplicative factor $S$ for the number of solver steps. For single-time-step prediction, RNNs and LSTMs require $\mathcal{O}(k)$, whereas Attention and NAC require $\mathcal{O}(nk)$ when recomputing attention over the full sequence.
Table 3: Sequence and time-step prediction complexity. $n$ is the sequence length and $k$ is the hidden/model dimension.
@@ C.3 Efficiency and Complexity
NAC-Exact $\mathcal{O}(n^{2}k)$ $\mathcal{O}(nk)$
NAC-Euler $\mathcal{O}(n^{2}k)$ $\mathcal{O}(nk)$
@@ D.1 Related Works
The brief description for related works is divided into four subcategories.
@@
DT-RNNs: RNN (rumelhart1985learning) captures sequential dependencies in time-series data by updating a hidden state from the current observation and the previous state. LSTM (hochreiter1997long) extends RNNs with input, output, and forget gates, allowing the network to maintain and update long-term memory, which improves modeling of long-term dependencies in time-series sequences. GRU (cho2014learning) simplifies the LSTM architecture by combining the forget and input gates into a single update gate, allowing efficient modeling of long-term dependencies in time-series sequences.
@@
CT-RNNs: CT-RNN (rubanova2019latent) model temporal dynamics using differential equations, enabling hidden states to evolve continuously over time in response to inputs, which is particularly useful for irregularly sampled time-series data. PhasedLSTM (neil2016phased) introduces a time gate that updates hidden states according to a rhythmic schedule, enabling efficient modeling of asynchronous or irregularly sampled time-series. GRU-ODE (de2019gru) extends the GRU to continuous time, evolving hidden states via ODEs to handle sequences with non-uniform time intervals. mmRNN (lechner2022mixed) combines short-term and long-term memory units to capture both fast-changing and slowly evolving patterns in sequential data. LTC (hasani2021liquid) use neurons with learnable, input-dependent time constants to adapt the speed of dynamics and capture complex temporal patterns in continuous-time data. CfC (hasani2022closed) approximate LTC dynamics analytically, providing efficient continuous-time modeling without relying on numerical ODE solvers.
@@
DT-Attentions: Attention (vaswani2017attention) computes attention weights by measuring similarity between queries and keys, scaling the results, and applying softmax to weigh time-step contributions. Multi-Head Attention (vaswani2017attention) applies multiple parallel scaled dot-product attention mechanisms, capturing different types of temporal dependencies simultaneously for complex time-series modeling.
@@
CT-Attentions: mTAN (shukla2021multi) learns continuous-time embeddings and uses time-based attention to interpolate irregular observations into a fixed-length representation for downstream encoder-decoder modeling. CTA (chien2021continuous) generalizes discrete-time attention to continuous-time by representing hidden states, context vectors, and attention scores as functions whose dynamics are modeled via neural networks and integrated using ODE solvers for irregular sequences. ODEFormer(d2023odeformer) trains a sequence-to-sequence transformer on synthetic trajectories to directly output a symbolic ODE system from noisy, irregular time-series data. ContiFormer (chen2023contiformer) builds a continuous-time Transformer by pairing ODE-defined latent trajectories with a time-aware attention mechanism to model dynamic relationships in irregular time-series data.
@@ D.2 Ablations Details
The brief descriptions of variants and ablation are also divided into four subcategories:
Top-K Ablations: NAC-2k uses Top-K=2 to compute the logits and NAC-32k uses Top-K=32. All variants use the exact computation mode with 50% sparsity.
Sparsity Ablations: NAC-02s uses 20% sparsity to compute the logits and NAC-09s uses 90%. NAC-PW employs full pairwise (non-sparse) concatenation for input curation. NAC-FC replaces the sparse NCP gating mechanism with a simple fully connected layer. All variants use the exact computation mode with Top-K=8.
Modes variants: NAC-Euler computes attention logits using the explicit Euler integration method. NAC-Steady derives attention logits from the steady-state solution of the exact formulation. NAC-Exact/05s/8k computes attention logits using the closed-form exact solution. It also overlaps with other ablations, so we combined it into a single one. All modes use Top-K=8, 50% sparsity and $\delta_{t}$=1.0. The sensitivity of NAC to $\delta_{t}$ is visualized in Figure 4
Figure 4: Effect of $\delta_{t}$ on output of NAC.
@@ D.3.1 Event-based MNIST
Preprocessing: We follow the preprocessing pipeline described in (lechner2022mixed), which proceeds as follows. First, a threshold is applied to convert the 8-bit pixel values into binary values, with 128 as the threshold on a scale from 0 (minimum intensity) to 255 (maximum intensity). Second, each $28\times 28$ image is reshaped into a one-dimensional time series of length 784. Third, the binary time series is encoded in an event-based format, eliminating consecutive occurrences of the same value; for example, the sequence $[1,1,1,1]$ is transformed into $(1,t=4)$. This encoding introduces a temporal dimension and compresses the sequences from 784 to an average of 53 time steps. Finally, to facilitate efficient batching and training, each sequence is padded to a fixed length of 256, and the time dimension is normalized such that each symbol corresponds to one unit of time. The resulting dataset defines a per-sequence classification problem on irregularly sampled time series.
@@ D.3.2 Person Activity Recognition (PAR)
Dataset Explanation and Curation: We used the Localized Person Activity Recognition dataset provided by UC Irvine (localization_data_for_person_activity_196). The dataset comprises 25 recordings of human participants performing different physical activities. The eleven possible activities are “walking,” “falling,” “lying down,” “lying,” “sitting down,” “sitting,” “standing up from lying,” “on all fours,” “sitting on the ground,” “standing up from sitting,” and “standing up from sitting on the ground.” The objective of this experiment is to recognize the participant’s activity from inertial sensors, formulating the task as a per-time-step classification problem. The input data consist of sensor readings from four inertial measurement units placed on participants’ arms and feet. While the sensors are sampled at a fixed interval of 211 ms, recordings exhibit different phase shifts and are thus treated as irregularly sampled time series.
Preprocessing: We first separated each participant’s recordings based on sequence identity and calculated elapsed time in seconds using the sampling period. To mitigate class imbalance, we removed excess samples from overrepresented classes to match the size of the smallest class. Subsequently, the data were normalized using a standard scaler. Finally, the dataset was split into a 90:10 ratio for training and testing.
Neural Network Architecture: Following the approach in Section D.3.1, we developed an end-to-end hybrid neural network combining convolutional heads with NAC or other baselines. Hyperparameter details are summarized in Table 4.
@@ D.3.3 Autonomous Vehicle
Dataset Explanation and Curation: We followed the data collection methodology described in (razzaq2023neural). For OpenAI-CarRacing, a PPO-trained agent (5M timesteps) was used to record 20 episodes, yielding approximately 48,174 RGB images of size $92\times 92\times 3$ with corresponding action labels across five discrete actions (no-act, move left, forward, move right, stop). The dataset was split with 10% reserved for testing and the remaining 90% for training. For the Udacity simulator, we manually controlled the vehicle for 50 minutes, producing 15647 RGB images of size $320\times 160\times 3$, captured from three camera streams (left, center, right) along with their corresponding continuous steering values. This dataset was split into 20% testing and 80% training.
@@
Neural Network Architecture: For OpenAI-CarRacing, we modified the neural network architecture proposed in (razzaq2023neural), which combines compact CNN layers for spatial feature extraction with LNNs to capture temporal dynamics. In our implementation, the LNN layers were replaced with NAC and its comparable alternatives for fair evaluation. Full hyperparameter configurations are provided in Table 4. For the Udacity simulator, we modified the network proposed in (bojarski2016end) by replacing three latent MLP layers with NAC and its counterparts. Full hyperparameters for this configuration are summarized in Table 4.
Saliency Maps: A saliency map visualizes the regions of the input that a model attends to when making decisions. Figure 5 shows the saliency maps for the OpenAI CarRacing environment. We observe that only NAC (Steady, Euler, and Steady) maintains focus on the road’s horizon, while other models either focus on the sides or remain largely unresponsive to the task. Figure 6 also presents the saliency maps for the Udacity Simulator. In this case, NAC-Exact, CTA produces the most accurate visual maps, maintaining attention on the road’s horizon, followed by ContiFormer and mTAN, which achieve comparable performance. Attention and PhasedLSTM also generate reasonable saliency maps, although their focus is more dispersed across the scene. In contrast, other models either fail to identify relevant regions, producing blurry maps, or focus solely on one side of the road. These results demonstrate the NAC’s ability to understand the underlying task.
@@
Algorithm 3 Time‒frequency Representation Algorithm
0:  windowed signal $I_{w}$, critical frequency $f_{c}$, operating frequency $f_{o}$, sampling period $T_{sampling}$, windowed physical constraints ($t_{w},T_{w}$)
  $a_{min}=\frac{f_{c}}{f_{max}\cdot T_{sampling}},\quad a_{{max}}=\frac{f_{c}}{f_{\text{min}}\cdot T_{sampling}}$,
  $a_{scale}\in[a_{min},a_{max}]$
  $I_{{TFR}}\leftarrow\{\}$
  for ($i_{w},t_{n},T_{n}$) in ($I_{w},t_{w},T_{w}$): do
  Wavelets: $\Gamma_{iw}(a,b)=\int_{-\infty}^{\infty}i_{w}\psi^{*}\left(\frac{t-b}{a}\right)dt.$
  Energy: $E=\sum_{m=1}^{M}\left|\Gamma_{iw}(a,b)\right|^{2}$
  Dominant frequency: $f_{d}=a_{\text{scale}}\left[\arg\max(E)\right]$.
  Entropy: $h=-\sum_{i=m}^{M}P(i_{w}(t))\log P(i_{w}(t))$.
  Kurtosis: $K=\frac{\mathbb{E}[(i_{w}(t)-\mu)^{4}]}{\sigma^{4}}$.
  Skewness: $s_{k}=\frac{\mathbb{E}[(i_{w}(t)-\mu)^{3}]}{\sigma^{3}}$.
  mean: $\mu=\frac{1}{M}\sum_{m=1}^{M}i_{w}(m)$.
  standard deviation: $\sigma=\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left(i_{w}(m)-\mu\right)^{2}}$.
  $X_{n}\leftarrow[\log(E),f_{d},h,K,sk,\mu,\sigma]$
  end for
  return $I_{TFR}=Concat(X_{1},X_{2}\ldots X_{N_{s}},t_{n},T_{n})$
Dataset Explanation and Curation: PRONOSTIA dataset is a widely recognized benchmark dataset in the field of condition monitoring and degradation estimation of rolling-element bearings. Nectiux et al.(nectoux2012pronostia) developed this dataset as part of the PRONOSTIA experimental platform. The dataset comprises 16 complete run-to-failure experiments performed under accelerated wear conditions across three different operating settings: 1800 rpm with 4 kN radial load, 1650 rpm with a 4.2 kN load, and 1500 rpm with a 5 kN load, all at a frequency of 100 Hz. Vibration data were recorded using accelerometers placed along the horizontal and vertical axes, which were sampled at 25.6 kHz. Additionally, temperature readings were collected at a sampling rate of 10 Hz. The data distributions for training and testing are provided in Table 5.
Table 5: Data distributions of PRONOSTIA, XJTU-SY, and HUST datasets.
Dataset Condition Frequency Radial Load Speed Train Test
PRONOSTIA Condition 1 100 Hz 4 kN 1800 rpm $4\sim 7$ $1\sim 3$
Condition 2 100 Hz 4.2 kN 1650 rpm - $1\sim 7$
Condition 3 100 Hz 5 kN 1500 rpm - $1\sim 3$
XJTU-SY Condition 1 35 Hz 4 kN 2100 rpm - $1\sim 5$
Condition 2 37.5 Hz 4.2 kN 2250 rpm - $1\sim 5$
Condition 3 40 Hz 5 kN 2400 rpm - $1\sim 5$
HUST Condition 1 - 0 W - - $1\sim 5$
Condition 2 - 200 W - - $1\sim 5$
Condition 3 - 400 W - - $1\sim 5$
Note: The PRONOSTIA dataset is utilized for training and generalization testing, while the XJTU-SY and HUST datasets are employed to evaluate cross-validation testing. (-) values are either not available or not utilized.
@@
Preprocess: Condition monitoring data comprises 1D non-stationary vibrational signals collected from multiple sensors. To extract meaningful information, these signals must be transformed into features that possess meaningful physical interpretability. We utilized the preprocessing proposed in (razzaq2025carle) and labels are generated according to (razzaq2025developingdistanceawareuncertaintyquantification). Initially, the signal is segmented into small, rectangularized vectors using a windowing technique ($w$), enabling better localization of transient characteristics. The continuous wavelet transform (CWT) (aguiar2014continuous) with the Morlet wavelet (lin2000feature) as the mother wavelet is then applied to obtain a time-frequency representation (TFR). The CWT is defined as $\Gamma(a,b)=\int_{-\infty}^{\infty}x_{w}(t)\frac{1}{\sqrt{a}}\psi^{*}\left(\frac{t-b}{a}\right)\,dt$, where $a$ and $b$ denote the scale and translation parameters, respectively, and $\psi$ is the Morlet wavelet function. From the resulting TFR, a compact set of statistical and domain-specific features is extracted to characterize the operational condition of the bearing. The complete feature extraction procedure is described in Algorithm 3.
@@
Evaluation Metric: Score is a metric specifically designed for RUL estimation in the IEEE PHM (nectoux2012pronostia) to score the estimates. The scoring function is asymmetric and penalizes overestimations more heavily than early predictions. This reflects practical considerations, as late maintenance prediction can lead to unexpected failures with more severe consequences than early intervention can.

### v3

@@ A.2 Attention Mechanism
    Normalize the logits to obtain attention weights and compute the output:
@@ Appendix B Proofs
B.1 Analyzing Closed-form Solution
B.1.1 Deriving Closed-form (Exact) Solution
Although $\phi$ and $\omega_{\tau}$ are nonlinear functions of the input $\mathbf{u}=[\mathbf{q};\mathbf{k}]$, we derive a closed-form solution by treating them as locally constant over the small integration interval for each query–key pair based on frozen-coefficient approximation (john1952integration). Rewrite Eqn. 1 as
@@ B.1.1 Deriving Closed-form (Exact) Solution
The integrating factor is
@@
Recognizing the left-hand side as the derivative of $e^{\omega_{\tau}t}a_{t}$:
@@
Compute the integral:
@@
Set $a^{*}:=\dfrac{\phi}{\omega_{\tau}}$. Then $a_{t}=a^{*}+(a_{0}-a^{*})e^{-\omega_{\tau}t}$, is proved.
B.1.2 Frozen Coefficient Error Analysis
In this subsection, we analyze the resulting error due to the frozen approximation with respect to the time-varying behavior of $\phi$ and $\omega_{\tau}$, both mathematically and empirically.
Theorem 3 (Frozen-Coefficient Approximation Error).
Consider the attention logit ODE
$\frac{da(t)}{dt}=-\omega_{\tau}(t)a(t)+\phi(t),$ (33)
with time-varying coefficients $\phi(t)$ and $\omega_{\tau}(t)>0$. Let $a_{\text{varying}}(t)$ denote the solution with varying coefficients (numerically integrated) and $a_{\text{frozen}}(t)$ denote the exact closed-form solution obtained by freezing the coefficients at their initial values $\phi_{0}=\phi(0)$, $\omega_{\tau 0}=\omega_{\tau}(0)$ over an integration window [0, T]. Assume that $\phi(t)$ and $\omega_{\tau}(t)$ are Lipschitz continuous in t with constants $L_{\phi}$ and $L_{\omega}$ (i.e., $|\phi(t)-\phi(s)|\leq L_{\phi}|t-s|$, $|\omega_{\tau}(t)-\omega_{\tau}(s)|\leq L_{\omega}|t-s|$). Define $C=|a_{0}|+|\phi_{0}|/\omega_{\tau 0}$. Then, for any $t\in[0,T]$, the error $\varepsilon(t)=a_{\text{varying}}(t)-a_{\text{frozen}}(t)$ satisfies
$|\varepsilon(t)|\leq(L_{\phi}+CL_{\omega})t^{2}e^{L_{\omega}t},$ (34)
For small $L_{\omega}t\ll 1$, this simplifies to $\mathcal{O}(t^{2})$.
Proof.
The frozen-coefficient solution satisfies
$\frac{da_{\text{frozen}}}{dt}=-\omega_{\tau 0}a_{\text{frozen}}+\phi_{0},\quad a_{\text{frozen}}(0)=a(0),$ (35)
Subtracting Eqn. 35 from Eqn. 33 yields the error dynamics
$\frac{d\varepsilon}{dt}=-\omega_{\tau 0}\varepsilon+\delta\phi(t)+\delta\omega_{\tau}(t)a_{\text{varying}}(t),$ (36)
where $\delta\phi(t)=\phi(t)-\phi_{0}$ and $\delta\omega_{\tau}(t)=\omega_{\tau 0}-\omega_{\tau}(t)$. By Lipschitz, $|\delta\phi(t)|\leq L_{\phi}t$ and $|\delta\omega_{\tau}(t)|\leq L_{\omega}t$. Using the integrating factor $e^{\omega_{\tau 0}t}$,
$\frac{d}{dt}\left(e^{\omega_{\tau 0}t}\varepsilon(t)\right)=e^{\omega_{\tau 0}t}\left(\delta\phi(t)+\delta\omega_{\tau}(t)a_{\text{varying}}(t)\right),$ (37)
Integrating from 0 to $T$ and noting that $\varepsilon(0)=0$,
$\varepsilon(t)=\int_{0}^{T}e^{-\omega_{\tau 0}(t-s)}\left(\delta\phi(s)+\delta\omega_{\tau}(s)a_{\text{varying}}(s)\right)ds,$ (38)
The frozen solution from Eqn. 32 is
$a_{\text{frozen}}(s)=a_{0}e^{-\omega_{\tau 0}s}+\frac{\phi_{0}}{\omega_{\tau 0}}\left(1-e^{-\omega_{\tau 0}s}\right),$ (39)
so that $|a_{\text{frozen}}(s)|\leq C$. Hence $|a_{\text{varying}}(s)|\leq C+|\varepsilon(s)|$. Taking absolute values in Eqn. 38 and using Lipschitz bounds,
$|\varepsilon(t)|\leq\int_{0}^{T}e^{-\omega_{\tau 0}(t-s)}\left(L_{\phi}s+L_{\omega}s(C+|\varepsilon(s)|)\right)ds,$ (40)
Since $e^{-\omega_{\tau 0}(t-s)}\leq 1$,
$|\varepsilon(t)|\leq(L_{\phi}+CL_{\omega})t^{2}+L_{\omega}\int_{0}^{T}|\varepsilon(s)|ds.$ (41)
Applying Gronwall’s inequality to $|\varepsilon(t)|$ yields
$|\varepsilon(t)|\leq(L_{\phi}+CL_{\omega})t^{2}e^{L_{\omega}t},$ (42)
which is Eqn. 34. Expanding the exponential for $L_{\omega}t\ll 1$ gives the $\mathcal{O}(t^{2})$ bound. ∎
Empirical Analysis: We now empirically analyze the error introduced by freezing $\phi(t)$ and $\omega_{\tau}(t)$. We evaluate the closed-form frozen solution against continuously varying coefficients. Empirical analysis across four regimes: (i) slow variation (gradual linear change); (ii) fast variation (exponential growth); (iii) step changes (abrupt discontinuities); and (iv) oscillatory behavior (periodic modulation) shows that the frozen approximation maintains bounded error even under substantial parameter variation. Table 2 summarizes the empirical results. For typical slow-varying irregular sequences, with $\phi$ varying by 0.5% and $\omega_{\tau}$ varying by 0.26%, the frozen solution exhibits minimal error, with a mean of 1.87%, and a maximum of 3.05%. Even under more challenging regimes, the error remains predictable and bounded, reaching 6.41% for fast variation, 6.53% for step changes, and 7.81% for oscillatory behavior. An illustration of oscillatory evolution is shown in Figure 5.
Interpretation: Given that the variation rates in real-world irregular sequences rarely exceed 10%, these results confirm the adequacy of the frozen approximation for practical deployment. We therefore recommend Exact as the default computation mode for NAC while preserving adaptive Euler integration for edge cases with exceptionally rapid dynamics.
Figure 5: Visualization of oscillatory evolution of NAC.
Table 2: Empirical error analysis of frozen-coefficient approximation across variation regimes.
Variation mean (%) max (%) $\phi$ (%) $\omega_{\tau}$ (%)
Slow 1.87 3.05 0.50 0.26
Fast 6.41 6.50 2.63 1.58
Step 6.53 7.20 3.51 2.11
Oscillatory 7.81 6.90 8.28 2.10
Average 5.65 5.18 3.73 1.51
@@ C.1.1 Trajectory sensitivities for Closed-form formulation
C.2 Efficiency and Complexity
Table 3 summarizes the computational complexity of different sequence models. For sequence prediction over length $n$ with hidden dimension $k$, RNNs scale linearly, $\mathcal{O}(nk)$, while Attention scale quadratically, $\mathcal{O}(n^{2}k)$. ODE-based models, such as LNNs, incur an additional multiplicative factor $S$ for the number of solver steps. NAC scale $\mathcal{O}(n\sqrt{n}k)$. For single-time-step prediction, RNNs, LSTMs, and NAC require $\mathcal{O}(k)$, whereas Attention and NAC require $\mathcal{O}(nk)$ when recomputing attention over the full sequence.
C.3 Gradient-Based Training
@@
Table 3: Sequence and time-step prediction complexity. $n$ is the sequence and $k$ is the hidden/model dimension.
@@
NAC-Exact $\mathcal{O}(n\sqrt{n}k)$ $\mathcal{O}(nk)$
@@ D.1 Related Works
The brief descriptions of related works are divided into four subcategories.
@@
DT-RNNs: RNN (rumelhart1985learning) captures sequential dependencies in time-series data by updating a hidden state from the current observation and the previous state. LSTM (hochreiter1997long) extends RNNs with input, output, and forget gates, allowing the network to maintain and update long-term memory, which improves the modeling of long-term dependencies in time-series sequences. GRU (cho2014learning) simplifies the LSTM architecture by combining the forget and input gates into a single update gate, allowing efficient modeling of long-term dependencies in time-series sequences.
@@
CT-RNNs: CT-RNN (rubanova2019latent) models temporal dynamics using differential equations, enabling hidden states to evolve continuously over time in response to inputs, which is particularly useful for irregularly sampled time-series data. PhasedLSTM (neil2016phased) introduces a time gate that updates hidden states according to a rhythmic schedule, enabling efficient modeling of asynchronous or irregularly sampled time-series. GRU-ODE (de2019gru) extends the GRU to continuous time, evolving hidden states via ODEs to handle sequences with nonuniform time intervals. mmRNN (lechner2022mixed) combines short-term and long-term memory units to capture both fast-changing and slowly evolving patterns in sequential data. LTC (hasani2021liquid) uses neurons with learnable, input-dependent time constants to adapt the speed of dynamics and capture complex temporal patterns in continuous-time data. CfC (hasani2022closed) approximate LTC dynamics analytically, providing efficient continuous-time modeling without relying on numerical ODE solvers. DeepState (rangapuram2018deep) combines RNNs with linear state-space models (SSMs), allowing the RNN to produce time-varying SSM parameters to model CT dynamics. S4 (gu2021efficiently) parametrizes SSMs with a structured low-rank correction to the state matrix, enabling efficient convolution-based computation and effective modeling of long-range dependencies.
@@
DT-Attentions: Attention (vaswani2017attention) computes attention weights by measuring the similarity between queries and keys, scaling the results, and applying Softmax to weigh time-step contributions. Multihead attention (vaswani2017attention) applies multiple parallel scaled dot-product attention mechanisms, capturing different types of temporal dependencies simultaneously for complex time-series modeling. Linear Attention (zhang2021sparse) replaces the softmax activation with a ReLU function to induce sparsity and enable heads to selectively switch off for certain queries. Performer (choromanski2020rethinking) approximates Softmax attention with linear-time FAVOR+, enabling efficient and scalable Attention without losing theoretical guarantees.
@@
CT-Attentions: mTAN (shukla2021multi) learns continuous-time embeddings and uses time-based attention to interpolate irregular observations into a fixed-length representation for downstream encoder-decoder modeling. CTA (chien2021continuous) generalizes discrete-time attention to continuous-time by representing hidden states, context vectors, and attention scores as functions whose dynamics are modeled via neural networks and integrated using ODE solvers for irregular sequences. ODEFormer(d2023odeformer) trains a sequence-to-sequence transformer on synthetic trajectories to directly output a symbolic ODE system from noisy, irregular time-series data. ContiFormer (chen2023contiformer) builds a CT-Transformer by pairing ODE-defined latent trajectories with a time-aware attention mechanism to model dynamic relationships in irregular time-series data.
@@ D.2 Ablations Details
The brief descriptions of variants and ablation are divided into five subcategories:
Top-K Ablations: NAC-2k uses Top-K=2 to compute the logits and NAC-32k uses Top-K=32. NAC-PW uses fully pairwise (no sparsity) concatenation. All variants use the exact computation mode with 50% sparsity.
NAC-Sparsity Ablations: NAC-02s uses 20% sparsity to compute the logits, and NAC-09s uses 90%. All variants use the exact computation mode with Top-K=8 and $\delta_{t}$=1.0.
NAC with isolated NCPs Ablations: To isolate the effect of the NCP gating mechanism, we ablate it using MLP layers with matched sparsity. NAC-FC replaces the gating module with a standard fully connected layer. NAC-09sFC applies 90% sparsity, NAC-05s applies 50% sparsity, and NAC-02s applies 20% sparsity. All variants use the exact computation mode with Top-$K=8$ and $\delta_{t}$=1.0.
NAC with unconstrained $\phi$: Applying a sigmoidal nonlinearity to $\phi$ yields the logit to only positive values. To allow negative logits, we tested two variants: (i) NAC-$\phi_{\text{linear}}$ with no activation, and (ii) NAC-$\phi_{\text{tanh}}$ with a tanh activation.
Modes Ablations: NAC-Euler computes attention logits using the explicit Euler integration method. NAC-Steady derives attention logits from the steady-state solution of the exact formulation. NAC-Exact/05s/8k computes attention logits using the closed-form exact solution. It also overlaps with other ablations, so we combined it into a single one. All modes use Top-K=8, 50% sparsity and $\delta_{t}$=1.0.
@@ D.3.1 Event-based MNIST
Preprocessing: We follow the preprocessing pipeline described in (lechner2022mixed), which proceeds as follows. First, a threshold is applied to convert the 8-bit pixel values into binary values, with 128 as the threshold on a scale from 0 (minimum intensity) to 255 (maximum intensity). Second, each $28\times 28$ image is reshaped into a one-dimensional time-series of length 784. Third, the binary time-series is encoded in an event-based format, eliminating consecutive occurrences of the same value; for example, the sequence $[1,1,1,1]$ is transformed into $(1,t=4)$. This encoding introduces a temporal dimension and compresses the sequences from 784 to an average of 53 time steps. Finally, to facilitate efficient batching and training, each sequence is padded to a fixed length of 256, and the time dimension is normalized such that each symbol corresponds to one unit of time. The resulting dataset defines a per-sequence classification problem on irregularly sampled time-series.
@@ D.3.2 Person Activity Recognition (PAR)
Dataset Explanation and Curation: We used the Localized Person Activity Recognition dataset provided by UC Irvine (localization_data_for_person_activity_196). The dataset comprises 25 recordings of human participants performing different physical activities. The eleven possible activities are “walking,” “falling,” “lying down,” “lying,” “sitting down,” “sitting,” “standing up from lying,” “on all fours,” “sitting on the ground,” “standing up from sitting,” and “standing up from sitting on the ground.” The objective of this experiment is to recognize the participant’s activity from inertial sensors, formulating the task as a per-time-step classification problem. The input data consist of sensor readings from four inertial measurement units placed on participants’ arms and feet. While the sensors are sampled at a fixed interval of 211 ms, recordings exhibit different phase shifts and are thus treated as irregularly sampled time-series.
Preprocessing: We first separated each participant’s recordings based on sequence identity and calculated elapsed time in seconds using the sampling period. To mitigate class imbalance, we removed excess samples from overrepresented classes to match the size of the smallest class. Subsequently, the data were normalized using a standard scaler. Finally, the dataset was split into 9:1 ratio for training and testing.
Neural Network Architecture: Following the approach in Section D.3.1, we developed an end-to-end neural network combining convolutional heads with NAC or other baselines. Hyperparameter details are summarized in Table 4.
@@ D.3.3 Autonomous Vehicle
Dataset Explanation and Curation: We followed the data collection methodology described in (razzaq2023neural). For OpenAI-CarRacing, a PPO-trained agent (5M timesteps) was used to record 20 episodes, yielding approximately 48174 RGB images of size $92\times 92\times 3$ with corresponding action labels across five discrete actions (no-act, move left, forward, move right, stop). The dataset was split with 10% reserved for testing and the remaining 90% for training. For the Udacity simulator, we manually controlled the vehicle for 50 minutes, producing 15647 RGB images of size $320\times 160\times 3$, captured from three camera streams (left, center, right) along with their corresponding continuous steering values. This dataset was split into 20% testing and 80% training.
@@
Neural Network Architecture: For OpenAI-CarRacing, we modified the neural network architecture proposed in (razzaq2023neural), which combines compact CNN layers for spatial feature extraction with LNNs to capture temporal dynamics. In our implementation, the LNN layers were replaced with NAC and its comparable alternatives for fair evaluation. Full hyperparameter configurations are provided in Table 4. For the Udacity simulator, we modified the network proposed in (bojarski2016end) by replacing three latent MLP layers with NAC and its counterparts. Full hyperparameters are summarized in Table 4.
Saliency Maps: A saliency map visualizes the regions of the input that a model attends to when making decisions. Figure 6 shows the saliency maps for the OpenAI CarRacing environment. We observe that only NAC (Steady, Euler, and Steady) maintains focus on the road’s horizon, while other models either focus on the sides or remain largely unresponsive to the task. Figure 7 also presents the saliency maps for the Udacity Simulator. In this case, NAC-Exact, CTA produce the most accurate visual maps, maintaining attention on the road’s horizon, followed by ContiFormer and mTAN, which achieve comparable performance. MHA, Linear Attention, Performer, and PhasedLSTM also generate reasonable saliency maps, although their focus is more dispersed across the scene. In contrast, other models either fail to identify relevant regions, producing blurry maps, or focus solely on one side of the road. These results demonstrate that NAC exhibits interpretable attention patterns in the underlying task.
@@
Dataset Explanation and Curation: PRONOSTIA dataset is a widely recognized benchmark dataset in the field of condition monitoring and degradation estimation of rolling-element bearings. Nectoux et al.(nectoux2012pronostia) developed this dataset as part of the PRONOSTIA experimental platform. The dataset comprises 16 complete run-to-failure experiments performed under accelerated wear conditions across three different operating settings: 1800 rpm with 4 kN radial load, 1650 rpm with a 4.2 kN load, and 1500 rpm with a 5 kN load, all at a frequency of 100 Hz. Vibration data were recorded using accelerometers placed along the horizontal and vertical axes, which were sampled at 25.6 kHz. Additionally, temperature readings were collected at a sampling rate of 10 Hz.
@@
Preprocessing: Condition monitoring data comprises 1D non-stationary vibrational signals collected from multiple sensors. To extract meaningful information, these signals must be transformed into features that possess meaningful physical interpretability. We utilized the preprocessing proposed in (razzaq2025carle), and physically plausible labels are generated according to (razzaq2025developingdistanceawareuncertaintyquantification). Initially, the signal is segmented into small, rectangularized vectors using a windowing technique, enabling better localization of transient characteristics. The continuous wavelet transform (CWT) with the Morlet wavelet as the mother wavelet is then applied to obtain a time-frequency representation (TFR). The CWT is defined as $\Gamma(a,b)=\int_{-\infty}^{\infty}x_{w}(t)\frac{1}{\sqrt{a}}\psi^{*}\left(\frac{t-b}{a}\right)\,dt$, where $a$ and $b$ denote the scale and translation parameters, respectively, and $\psi$ is the Morlet wavelet function. From the TFR, a compact set of statistical and domain-specific features is extracted to characterize the operational condition of the bearing.
@@
Evaluation Metric: Score is a metric specifically designed for RUL estimation in the IEEE PHM (nectoux2012pronostia) to score the estimates. The scoring function is asymmetric, penalizing overestimations more heavily than early predictions. This reflects practical considerations, as late maintenance prediction can lead to unexpected failures with more severe consequences than early intervention can.